Total Cost of Employee B. F. Goodson employs Eduardo Gonzales at a salary of $47,900 a year. Goodson is subject to employer Social Security taxes at a rate of 6.2% and Medicare taxes at a rate of 1.45% on Gonzales's salary. In addition, Goodson must pay SUTA tax at a rate of 5.9% and FUTA tax at a rate of 0.7% on the first $7,000 of Gonzales's salary. Compute the total cost to Goodson of employing Gonzales for the year. If required, round your answer to the nearest cent.

$_______
Answer #1
Salary expense 47900.00
Social Security tax 2969.80 =47900*6.2%
Medicare tax 694.55 =47900*1.45%
SUTA tax 413.00 =7000*5.9%
FUTA tax 49.00 =7000*0.7%
Total cost to Goodson of employing Gonzales 52026.35